Solve each equation on the interval 0 1. cos^2(x)-sin^2(x)+sin(x)=0
2. sin^2(x)= 6(cosx+1)
3. 2sin^2(x)= 3(1-cosx)
4. cos(x)=sin(x)
5. cos(x)+sin(x)=0
6. tan(x)=2sin(x)
7. 1+sin(x)=2cos^2(x)
8. sin^2(x)=2cos(x)+2
9. tan^2(x)= 3/2 sec(x)
10. csc^2(x)=cot(x)+1
11. sec^2(x)+tan(x)=0
12. sec(x)=tan(x)+cot(x)
13. sin(x)- [square root of 3]cos(x)= 1

Answer by Edwin McCravy(6938)

Solve each equation on the interval 0 

That's way too many problems to post!  I'll solve only 3,6, and 12. 
and give only the answers to the rest. BTW, 0 is a number, not an 
interval.  So I can't tell if you want them in
degrees or radians.  I gave them in radians. If you want degrees,
multiply by 180/pi.

1. cos^2(x)-sin^2(x)+sin(x)=0           pi/2, 7pi/6, 11pi/6
2. sin^2(x)= 6(cosx+1),                 pi
3. 2sin^2(x)= 3(1-cosx),                0, pi/3, 5pi/3 

                             2(1 - cos²x) = 3(1 - cosx)
                    2(1 - cosx)(1 + cosx) = 3(1 - cosx)
      2(1 - cosx)(1 + cosx) - 3(1 - cosx) = 0    
              (1 - cosx)[2(1 + cosx) - 3] = 0
                (1 - cosx)[2 + 2cosx - 3] = 0
                    (1 - cosx)(2cosx - 1) = 0 

setting first factor = 0

                                 1 - cosx = 0
                                    -cosx = -1
                                     cosx = 1
                                        x = 0 

setting second factor = 0

                                2cosx - 1 = 0
                                    2cosx = 1
                                     cosx = 1/2
                                        x = pi/3, 5pi/3 

4. cos(x)=sin(x),                       pi/4, 5pi/4 
5. cos(x)+sin(x)=0                     3pi/4, 7pi/4
6. tan(x)=2sin(x),                     0, pi/3, pi, 5pi/3 

                               sinx/cosx = 2sinx
                                    sinx = 2sinx·cosx
                       sinx - 2sinx·cosx = 0 
                         sinx(1 - 2cosx) = 0

setting first factor = 0

                                    sinx = 0
                                       x = 0, pi

setting second factor = 0   

                               1 - 2cosx = 0
                                  -2cosx = -1
                                    cosx = 1/2
                                       x = pi/3, 5pi/3

7. 1+sin(x)=2cos^2(x),                 pi/6, 5pi/6, 3pi/2
8. sin^2(x)=2cos(x)+2,                 pi
9. tan^2(x)= 3/2 sec(x),               pi/3, 5pi/3
10. csc^2(x)=cot(x)+1,                 pi/4, pi/2, 5pi/4, 3pi/2 
11. sec^2(x)+tan(x)=0,                 no solution
12. sec(x)=tan(x)+cot(x),              no solution

                                1/cosx = sinx/cosx + cosx/sinx

Multiply thru by LCD = cosx·sinx

                                  sinx = sin²x + cos²x
                                  sinx = 1
                                     x = pi/2

              However we must discard that solution
              because the original problem contains
              tanx, and tan(pi/2) is not defined.

13. sin(x)- [square root of 3]cos(x)= 1, pi/2, 7pi/6

Edwin